OpenStudy (anonymous):

simplify by removing factors of 1 n^2-1/(n+1)^2

OpenStudy (anonymous):

\[\frac{n^2-1}{(n+1)^2}=\frac{(n-1)(n+1)}{(n+1)^2}\]Cancel the common factor, what do you get?

OpenStudy (anonymous):